# Preference-Guided Reinforcement Learning for Efficient Exploration

### v1

@@ Abstract
In this paper, we investigate preference-based reinforcement learning (PbRL) that allows reinforcement learning (RL) agents to learn from human feedback. This is particularly valuable when defining a fine-grain reward function is not feasible. However, this approach is inefficient and impractical for promoting deep exploration in hard-exploration tasks with long horizons and sparse rewards. To tackle this issue, we introduce LOPE: Learning Online with trajectory Preference guidancE, an end-to-end preference-guided RL framework that enhances exploration efficiency in hard-exploration tasks. Our intuition is that LOPE directly adjusts the focus of online exploration by considering human feedback as guidance, avoiding learning a separate reward model from preferences. Specifically, LOPE includes a two-step sequential policy optimization process consisting of trust-region-based policy improvement and preference guidance steps. We reformulate preference guidance as a novel trajectory-wise state marginal matching problem that minimizes the maximum mean discrepancy distance between the preferred trajectories and the learned policy. Furthermore, we provide a theoretical analysis to characterize the performance improvement bound and evaluate the LOPE’s effectiveness. When assessed in various challenging hard-exploration environments, LOPE outperforms several state-of-the-art methods regarding convergence rate and overall performance. The code used in this study is available at https://github.com/buaawgj/LOPE.
Index Terms:
Deep reinforcement learning, hard exploration, trajectory preference, state marginal matching, two-step policy optimization.
@@ I Introduction
Deep reinforcement learning (DRL) has recently successfully solved challenging problems through trial and error [1, 2, 3]. However, many real-world tasks involve long horizons and poorly defined goals and specifying a suitable reward function in these tasks is difficult [4, 5]. Such hard-exploration tasks remain exceedingly challenging for DRL [6, 7]. It is difficult for agents to explore efficiently and obtain highly rewarded trajectories in such environments [6]. This realistic problem further imposes an extreme urgency for efficient exploration, and overcoming this drawback can considerably expand the possible impact of RL.
@@
An alternative to mitigate the sparsity of reward feedback is preference-based reinforcement learning (PbRL) [8, 9]. In PbRL, instead of directly receiving the instant reward information on each encountered state-action pair, the agent only obtains 1-bit preference feedback for each state-action pair or trajectory from a human overseer [10, 11]. To acquire a dense reward function, many prior PbRL methods employ specialized models to learn separate reward functions from human preferences and then train policies using these learned reward functions [12, 13]. Such a method suggests that the agent is instructed to act optimally indirectly. Furthermore, it increases model complexity and inexplainability and creates a potential information bottleneck.
@@
Most existing PbRL approaches do not consider scaling to hard-exploration tasks [13, 11]. Although learning a separate reward function does help in solving the sparse reward problem, continuously enforcing such preference-based rewards during the whole training phase cannot guarantee deep exploration in hard-exploration tasks. Only a few studies encourage exploration based on uncertainty in learned reward functions for PbRL algorithms [14]. However, we have no guarantee of achieving deep exploration with this method in hard-exploration tasks with long horizons and large state spaces. Therefore, scaling PbRL to hard-exploration tasks can fill this research gap of RL.
@@
In this study, we propose a novel online RL algorithm called Learning Online with trajectory Preference guidancE (LOPE), an end-to-end framework that jointly encourages exploration and learns the optimal control policy by regarding human preferences as guidance. Our intuition is that LOPE orients its policy according to human preferences and fully utilizes the distribution information of preferences. LOPE involves a two-step sequential policy optimization process, including trust-region-based policy improvement and preference guidance steps. Specifically, the first step is to employ a trust-region-based approach to generate a candidate policy. In the second step, we introduce a novel trajectory-based distance between policies based on the maximum mean discrepancy (MMD) and reformulate preference guidance as a novel trajectory-wise state marginal matching problem. We then show this trajectory-wise state marginal matching problem can be transformed into a policy-gradient algorithm with shaped rewards learned from human preferences. Furthermore, this approach enables us to provide theoretical performance guarantees for performance improvement while achieving outstanding performance in various benchmark tasks. Extensive experimental results illustrate the superior performance of LOPE against competitive baselines.
@@
Our contributions are summarized as follows: (1) LOPE: a simple, efficient, and end-to-end PbRL approach that addresses the exploration difficulty by regarding human preferences as guidance; (2) a novel preference-guided trajectory-wise state marginal matching optimization objective and a two-step optimization framework; (3) no requirement for additional neural networks; (4) analytical performance guarantees of LOPE.
@@
Figure 1: Illustration of our method. First, the agent interacts with the environment by the policy $\pi$ and collects a set $\mathcal{B}$ of trajectories $\{\tau^{1},\dots,\tau^{K}\}$. Then, we calculate the policy gradient based on environmental rewards and update policy parameters to increase the expected return and update the current policy toward human preferences. We send the set of trajectories $\mathcal{B}$ to the annotator for a comparison with trajectories in the preferred-trajectory set $\mathcal{P}$ and update $\mathcal{P}$.
@@ II-A Preference-based RL
Preference-based RL is viewed as human-in-the-loop reinforcement learning by learning from human feedback [15]. Overall, there are three main types of preference feedback in the recent PbRL work. The first category of human preference indicates the preferred states between different state pairs, where actions that are better than all available actions exist in the other state [15]. Second, action preferences are defined on the action space to determine the better action in a given state [16, 17, 18, 19, 8]. Lastly, trajectory preferences compare the human feedback between trajectory pairs and specify the preferred trajectories over the other ones [12, 10].
Trajectory preferences are the most popular form among the three types of human feedback and are the main focus of many researches [11]. Recently, some impressive work introduced some theoretical results about PbRL with trajectory feedback, which mainly concentrate on the tabular RL setting[20, 21, 22].  [11] propose an end-to-end offline preference-guided RL algorithm to avoid the potential information bottleneck. SURF infers the pseudo-labels of unlabeled samples based on the confidence of the preference predictor for reward learning [23]. However, these approaches often require learning a separate reward function corresponding to human feedback or an extra model to encode preference information, causing a higher computation burden and unstable training.
@@ II-B Learn from Demonstrations (LfD)
The objective of LfD is to improve the policy exploration ability of the agent and accelerate learning using the combination of RL with expert demonstration experience [24]. Recently, various work has focused on the research field of LfD. [25] uses expert demonstrations to design an auxiliary margin loss for a pre-trained deep Q-function. [26] shows that model-free DRL algorithms can scale to high-dimensional dexterous manipulation by behavior cloning from a few human expert demonstrations. AWAC accelerates learning of online RL by leveraging a prodigious amount of offline demonstration data with associated rewards [27]. LOGO exploits offline and imperfect demonstrations generated by a sub-optimal policy to achieve faster and more efficient online RL policy optimization [28]. Self-imitation learning (SIL) methods train the agent to imitate its past self-generated experience only when the return of the trajectory exceeds the value function of a certain threshold [29, 30, 31]. However, LfD methods place a high demand on high-quality demonstrations, and the performance of these methods is limited by the expert policy that generated the demonstration data.
@@ II-C RL with Once-per-episode Feedback
The goal of RL with once-per-episode feedback is to design a new paradigm to cope with the lack of reward functions in various realistic scenarios in which a reward is only revealed at the end of the episode. [32] propose to use an online linear bandit algorithm [33] to directly estimate the underlying Markov reward function and introduce a hybrid optimistic-Thompson Sampling method with a $\sqrt{K}$ regret. Meanwhile, [34] abandon the Markov reward assumption to provide an efficient algorithm based on the recent work [35], obtaining its regret bounds with the UCBVI algorithm [36].  [37] propose the PARTED algorithm that decomposes the trajectory-wise return into a reward sequence for every state-action pair in the trajectory and then performs a pessimistic value function update using the learned rewards. Unlike this, our method exploits human feedback to guide agents to explore environments efficiently and can be applied to more complicated scenes.
@@ III-A Reinforcement Learning
We formulate the reinforcement learning problem as a discrete-time Markov decision process (MDP) [38] described by a tuple $(\mathcal{S},\mathcal{A},r,P,\rho_{0},\gamma)$. Here, $\mathcal{S}$ and $\mathcal{A}$ denote the spaces of states and actions, respectively. $\gamma\in(0,1)$ is the discount factor. A stochastic policy $\pi(a_{t}|s_{t})$ defines an action probability distribution conditioned by $s_{t}$. At each time step $t$, given an action $a_{t}$ from $\pi$, the agent samples the next state $s_{t+1}$ from the transition probability distribution $P$, i.e., $s_{t+1}\sim P(s_{t+1}|s_{t},a_{t})$, and it receives a reward $r(s_{t},a_{t})$ determined by the reward function $r:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$. $\rho_{0}$ is the initial state distribution. The objective function is expressed as the expectation of discounted cumulative rewards:
@@ III-B PbRL with Trajectory Preferences
In this study, we consider a PbRL framework with trajectory preferences. Unlike the standard RL problem setting, where the agent can receive per-step proxy rewards, we do not assume such instantaneous rewards are accessible. Instead, the agent can only obtain trajectory preferences revealed by an annotator during the early training process. Given these trajectory preferences, suppose that there is a reward function $r_{\psi}(s_{t},a_{t})$ that coincides with these human preferences, the goal of PbRL algorithms is to learn a policy $\pi(a_{t}|s_{t})$ to maximize the expected (discounted) cumulative rewards $J_{r_{\psi}}$. To achieve this goal, vast amounts of previous work learn a reward function with these human preferences [12, 39, 40, 41], which is supposed to be consistent with the expert preferences. In contrast, our method directly the policy from these trajectory preferences.
@@ III-C State Marginal Matching
State marginal matching (SMM) has been recently studied as an alternative problem in RL [42, 43, 44]. Different from the RL objective to maximize the expected return, the SMM objective is to find a policy whose state marginal distribution $\rho_{\pi}(s)$ minimizes the divergence $D$ to a given target distribution $\rho^{\ast}(s)$:
@@
where $D$ is a divergence measure such as Kullback-Leibler (KL) divergence [42] and Jensen-Shannon (JS) divergence [45]. For the target distribution $\rho^{\ast}(s)$, [42] utilize a uniform distribution to improve exploration in the entire state space; some other approaches leverage imitation learning to match a small set of expert demonstrations [45, 43, 46].
@@ III-D Maximum Mean Discrepancy
The maximum mean discrepancy (MMD) is an integral probability metric that can be used to measure the difference (or similarity) between two different probability distributions [47, 48, 49]. Suppose that $p$ and $q$ are the probability distributions defined on a nonempty compact metric space $\mathbb{X}$. Let $x$ and $y$ be observations sampled independently from $p$ and $q$, respectively. Given a Reproducing Kernel Hilbert Space (RKHS) $\mathcal{H}$, the MMD metric is given as:
@@ IV Learning Online with Trajectory Preference Guidance
This section presents a novel online PbRL algorithm called Learning Online with trajectory Preference guidancE (LOPE). Our main idea is to consider trajectory preferences as guidance for direct policy optimization in hard-exploration tasks, avoiding the potential information bottleneck incurred by learning separate reward functions from these preferences. We first introduce a novel preference-guided trajectory-wise state marginal matching optimization objective by reformulating a constrained optimization problem. Then, we develop a two-step policy optimization framework by integrating a policy improvement step with an additional preference guidance step.
@@ IV-A Human-in-loop: Trajectory-Wise State Marginal Matching
Many previous works of PbRL explicitly learn reward functions coinciding with human preferences. Alternatively, we seek to perform direct policy optimization with a preference-labeled data set. Inspired by previous research [42, 43], we propose reforming preference guidance as a trajectory-wise state marginal matching problem. Specifically, we define the divergence $d(\tau,\upsilon)$ between state visitation distributions of trajectories $\tau$ and $\upsilon$. Then, based on $d(\cdot,\cdot)$, we obtain the following trajectory-wise state marginal matching objective between $\pi$ and $\pi_{b}$:
@@
Compared with previous studies of state marginal matching [42, 43], our approach emphasizes trajectory-level alignment. During training, many state-action pairs are concentrated near the starting point, and only on rare occasions can the agent visit novel regions. Hence, due to the low data proportion, these novel state-action pairs have little impact on policy optimization when the SMM objective only considers the divergence between different state visitation distributions. Our approach solves this problem by introducing a trajectory-wise SMM objective and drives the agent to achieve deep exploration.
Now the focus of the problem comes to the definition of $d(\cdot,\cdot)$. Inspired by previous research [48, 50], a trajectory can be considered a deterministic policy, and the distance between them can be defined as the MMD between their state visitation distributions. Therefore, $d(\cdot,\cdot)$ can be expressed as:
@@
$+\underset{s_{\upsilon},s_{\upsilon}^{\prime}\sim\rho_{\upsilon}}{\mathbb{E}}\left[k(s_{\upsilon},s_{\upsilon}^{\prime})\right],$
where $\rho_{\tau}$ and $\rho_{\upsilon}$ are the state visitation distributions of $\tau$ and $\upsilon$. To guarantee the stability of policy optimization, we further adopt the trust-region-based optimization method similar to TRPO [51]:
@@
[…]
Intuitively, this preference-guided objective function aids learning by orienting the agent’s policy toward human feedback. Specifically, our approach finds a new policy $\pi_{new}$ that minimizes the expected trajectory-wise MMD distance to $\pi_{b}$, while ensuring $\pi_{new}$ to lie inside the trust region around $\pi_{old}$. This preference-guided policy optimization is the key distinguishing LOPE from other PbRL approaches. This method equips LOPE with two unique advantages over other state-of-the-art algorithms.
First, many previous PbRL studies explicitly learn separate reward functions and then employ off-the-shelf RL algorithms to train a control policy. In contrast, LOPE considers preference-labeled trajectories as guidance and reformulates a novel constrained optimization problem in Eq. (6). In this manner, LOPE effectively avoids the potential information bottleneck incurred by conveying information from preferences to the policy via the scalar rewards.
Second, LOPE learns a policy matching the trajectory-wise state visitation distribution of preference-labeled experiences. Consequently, LOPE enables the agent to quickly expand the scope of exploration along the preference-labeled trajectories in $\mathcal{P}$ and overcome the challenging exploration difficulty in online RL settings. Fundamentally, the proposed approach allows us to explore the available information of preference data more comprehensively, such as position information, allowing more significant assistance to RL algorithms.

### v2

@@ Abstract
In this paper, we investigate preference-based reinforcement learning (PbRL), which enables reinforcement learning (RL) agents to learn from human feedback. This is particularly valuable when defining a fine-grain reward function is not feasible. However, this approach is inefficient and impractical for promoting deep exploration in hard-exploration tasks with long horizons and sparse rewards. To tackle this issue, we introduce LOPE: Learning Online with trajectory Preference guidancE, an end-to-end preference-guided RL framework that enhances exploration efficiency in hard-exploration tasks. Our intuition is that LOPE directly adjusts the focus of online exploration by considering human feedback as guidance, thereby avoiding the need to learn a separate reward model from preferences. Specifically, LOPE includes a two-step sequential policy optimization technique consisting of trust-region-based policy improvement and preference guidance steps. We reformulate preference guidance as a trajectory-wise state marginal matching problem that minimizes the maximum mean discrepancy distance between the preferred trajectories and the learned policy. Furthermore, we provide a theoretical analysis to characterize the performance improvement bound and evaluate the effectiveness of the LOPE. When assessed in various challenging hard-exploration environments, LOPE outperforms several state-of-the-art methods in terms of convergence rate and overall performance.
@@ I Introduction
Deep reinforcement learning (DRL) has recently successfully solved challenging problems through trial and error [mnih2015human, mnih2016asynchronous, silver2016mastering]. However, many real-world tasks involve long horizons and poorly defined goals and specifying a suitable reward function in these tasks is difficult [Jeon2020RewardrationalC, HadfieldMenell2017InverseRD]. Such hard-exploration tasks remain exceedingly challenging for DRL [yang2021exploration, guo2020memory]. It is difficult for agents to explore efficiently and obtain highly rewarded trajectories in such environments [yang2021exploration]. This realistic problem further imposes an extreme urgency for efficient exploration, and overcoming this drawback can considerably expand the possible impact of RL.
@@
An alternative approach to mitigate the sparsity of reward feedback is preference-based reinforcement learning (PbRL) [wirth2016model, abdelkareem2022advances]. In PbRL, instead of directly receiving the instant reward information on each encountered state-action pair, the agent only obtains 1-bit preference feedback for each state-action pair or trajectory from a human overseer [Chen2022HumanintheloopPE, kang2023beyond]. To acquire a dense reward function, many prior PbRL methods employ specialized models to learn separate reward functions from human preferences and then train policies using these learned reward functions [christiano2017deep, lee2021pebble]. Such a method suggests that the agent is instructed to act optimally indirectly. Furthermore, it increases model complexity and in-explainability, creating a potential information bottleneck.
@@
Most existing PbRL approaches do not consider scaling to hard-exploration tasks [lee2021pebble, kang2023beyond]. Although learning a separate reward function helps in solving the sparse reward problem, continuously enforcing such preference-based rewards throughout the entire training phase cannot guarantee deep exploration in hard-exploration tasks. Only a few studies encourage exploration based on uncertainty in learned reward functions for PbRL algorithms [liang2022reward]. However, we have no guarantee of achieving deep exploration with this method in hard-exploration tasks with long horizons and large state spaces. Therefore, scaling PbRL to hard-exploration tasks can fill this research gap in RL.
@@
In this study, we propose an online RL algorithm called Learning Online with trajectory Preference guidancE (LOPE). This end-to-end framework jointly encourages exploration and directly learns the optimal control policy by regarding human preferences as guidance. LOPE is designed as an end-to-end framework that directly updates policies based on human preferences and environmental feedback, without requiring an explicit reward model or a separate reward learning phase, as shown in Fig 1. Each training iteration consists of two tightly coupled components: a trust-region-based policy improvement step and a preference-guided exploration mechanism. These components are executed jointly in every iteration and together form a unified policy optimization process. Specifically, the first step involves employing a trust-region-based approach to generate a candidate policy based on environmental rewards. In the second step, we introduce a trajectory-based distance between policies based on the maximum mean discrepancy (MMD) and reformulate preference guidance as a trajectory-wise state marginal matching problem. We then show that this trajectory-wise state marginal matching problem can be transformed into a policy-gradient algorithm with shaped rewards learned from human preferences. Furthermore, this approach enables us to provide theoretical performance guarantees for performance improvement while achieving outstanding performance in various benchmark tasks. Extensive experimental results demonstrate the superior performance of LOPE compared to competitive baselines.
@@
Our contributions are summarized as follows: (1) LOPE: a simple, efficient, and end-to-end PbRL approach that addresses the exploration difficulty by regarding human preferences as guidance; (2) a preference-guided trajectory-wise state marginal matching optimization objective and a unified two-step optimization technique in each policy iteration; (3) no requirement for additional neural networks; (4) analytical performance guarantees of LOPE.
@@
Figure 1: Illustration of our method. Due to the hard exploration and sparse reward, the traditional policy improvement is inefficient. In this work, we propose an end-to-end preference-guided RL framework to achieve efficient exploration.
@@ II-A Preference-based RL
Preference-based RL is viewed as human-in-the-loop reinforcement learning, as it learns from human feedback [wirth2017survey]. Notably, recent studies have proposed using transformer-based models to better capture preference dynamics [kim2023preference], learned dynamics models to guide preference-efficient policy learning [liu2023efficient], and population-based mechanisms for improved exploration in preference space [driss2025pb]. Unlike our method, these approaches employ more complex structures to achieve preference modeling and efficient exploration. Overall, there are three main types of preference feedback in the recent PbRL work. The first category of human preference indicates the preferred states between different state pairs, where actions that are better than all available actions exist in the other state [wirth2017survey]. Second, action preferences are defined on the action space to determine the better action in a given state [Sugiyama2012PreferencelearningBI, Daniel2015ActiveRL, Asri2016ScorebasedIR, Arumugam2019DeepRL, wirth2016model]. Lastly, trajectory preferences compare the human feedback between trajectory pairs and specify the preferred trajectories over the other ones [christiano2017deep, Chen2022HumanintheloopPE].
Trajectory preferences are the most popular form among the three types of human feedback and are the main focus of many research [kang2023beyond]. Recently, some impressive work introduced some theoretical results about PbRL with trajectory feedback, which mainly concentrates on the tabular RL setting[Novoseller2019DuelingPS, Xu2020PreferencebasedRL, Pacchiano2021DuelingRR].  [kang2023beyond] proposes an end-to-end offline preference-guided RL algorithm to avoid the potential information bottleneck. SURF infers the pseudo-labels of unlabeled samples based on the confidence of the preference predictor for reward learning [parksurf]. However, these approaches often require learning a separate reward function corresponding to human feedback or an additional model to encode preference information, resulting in a higher computational burden and unstable training.
@@ II-B Learn from Demonstrations (LfD)
The objective of LfD is to enhance the policy exploration ability of the agent and accelerate learning by combining RL with expert demonstration experience [schaal1996learning]. Recently, various works have focused on the research field of LfD. [hester2018deep] uses expert demonstrations to design an auxiliary margin loss for a pre-trained deep Q-function. [rajeswaranlearning] shows that model-free DRL algorithms can scale to high-dimensional dexterous manipulation by behavior cloning from a few human expert demonstrations. AWAC accelerates learning of online RL by leveraging a prodigious amount of offline demonstration data with associated rewards [nair2020awac]. LOGO exploits offline and imperfect demonstrations generated by a sub-optimal policy to achieve faster and more efficient online RL policy optimization [rengarajanreinforcement]. Self-imitation learning (SIL) methods train the agent to imitate its past self-generated experience only when the return of the trajectory exceeds the value function of a certain threshold [oh2018self, gangwani2019learning, libardi2021guided]. However, LfD methods place a high demand on high-quality demonstrations, and the performance of these methods is limited by the expert policy that generated the demonstration data.
@@ II-C RL with Once-per-episode Feedback
The goal of RL with once-per-episode feedback is to design a new paradigm that addresses the lack of reward functions in various realistic scenarios where a reward is only revealed at the end of the episode. [efroni2021reinforcement] propose to use an online linear bandit algorithm [abbasi2011improved] to directly estimate the underlying Markov reward function and introduce a hybrid optimistic-Thompson Sampling method with a $\sqrt{K}$ regret. Meanwhile, [chatterji2021theory] abandons the Markov reward assumption to provide an efficient algorithm based on the recent work [russac2021self], obtaining its regret bounds with the UCBVI algorithm [azar2017minimax].  [xu2022provably] proposes the PARTED algorithm that decomposes the trajectory-wise return into a reward sequence for every state-action pair in the trajectory and then performs a pessimistic value function update using the learned rewards. Unlike this, our method utilizes human feedback to guide agents in exploring environments efficiently and can be applied to more complex scenes.
@@ III-A Reinforcement Learning
We formulate the reinforcement learning problem as a discrete-time Markov decision process (MDP) [sutton2018reinforcement] described by a tuple $(\mathcal{S},\mathcal{A},r,P,\rho_{0},\gamma)$. Here, $\mathcal{S}$ and $\mathcal{A}$ denote the spaces of states and actions, respectively. $\gamma\in(0,1)$ is the discount factor. A stochastic policy $\pi(a_{t}|s_{t})$ defines an action probability distribution conditioned by $s_{t}$. At each time step $t$, given an action $a_{t}$ from $\pi$, the agent samples the next state $s_{t+1}$ from the transition probability distribution $P$, i.e., $s_{t+1}\sim P(s_{t+1}|s_{t},a_{t})$, and it receives a reward $r(s_{t},a_{t})$ determined by the reward function $r:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$. $\rho_{0}$ is the initial state distribution. The objective function is expressed as the expectation of discounted cumulative rewards:
@@ III-B PbRL with Trajectory Preferences
In this study, we consider a PbRL framework with trajectory preferences. Unlike the standard RL problem setting, where the agent can receive per-step proxy rewards, we do not assume that such instantaneous rewards are accessible. Instead, the agent can only obtain trajectory preferences revealed by an annotator during the early training process. Given these trajectory preferences, suppose that there is a reward function $r_{\psi}(s_{t},a_{t})$ that coincides with these human preferences, the goal of PbRL algorithms is to learn a policy $\pi(a_{t}|s_{t})$ to maximize the expected (discounted) cumulative rewards $J_{r_{\psi}}$. To achieve this goal, vast amounts of previous work have learned a reward function that incorporates these human preferences [christiano2017deep, ibarz2018reward, stiennon2020learning, wu2021recursively], which is intended to be consistent with the expert preferences. In contrast, our method directly extracts the policy from these trajectory preferences.
@@ III-C State Marginal Matching
State marginal matching (SMM) has been recently studied as an alternative problem in RL [lee2019efficient, ghasemipour2020divergence, furuta2021generalized]. Different from the RL objective to maximize the expected return, the SMM objective is to find a policy whose state marginal distribution $\rho_{\pi}(s)$ minimizes the divergence $D$ to a given target distribution $\rho^{\ast}(s)$:
@@
where $D$ is a divergence measure such as Kullback-Leibler (KL) divergence [lee2019efficient] and Jensen-Shannon (JS) divergence [sharma2022state]. For the target distribution $\rho^{\ast}(s)$, [lee2019efficient] utilizes a uniform distribution to improve exploration in the entire state space; some other approaches leverage imitation learning to match a small set of expert demonstrations [sharma2022state, ghasemipour2020divergence, kostrikov2019imitation].
@@ III-D Maximum Mean Discrepancy
The maximum mean discrepancy (MMD) is an integral probability metric that can be used to measure the difference (or similarity) between two different probability distributions [gretton2012optimal, masood2019diversity, wang2024trajectory]. Suppose that $p$ and $q$ are the probability distributions defined on a nonempty compact metric space $\mathbb{X}$. Let $x$ and $y$ be observations sampled independently from $p$ and $q$, respectively. Given a Reproducing Kernel Hilbert Space (RKHS) $\mathcal{H}$, the MMD metric is given as:
@@ IV Learning Online with Trajectory Preference Guidance
This section presents a novel online PbRL algorithm called Learning Online with trajectory Preference guidancE (LOPE). Our primary objective is to utilize trajectory preferences as guidance for direct policy optimization in hard-exploration tasks, thereby avoiding the potential information bottleneck that arises from learning separate reward functions from these preferences. We first introduce a novel preference-guided trajectory-wise state marginal matching optimization objective by reformulating a constrained optimization problem. Then, we develop a two-step policy optimization framework by integrating a policy improvement step with an additional preference guidance step.
@@ IV-A Human-in-loop: Trajectory-Wise State Marginal Matching
Many previous works on PbRL have explicitly learned reward functions that coincide with human preferences. Alternatively, we aim to perform direct policy optimization using a preference-labeled dataset. Inspired by previous research [lee2019efficient, ghasemipour2020divergence], we propose reforming preference guidance as a trajectory-wise state marginal matching problem. Specifically, we define the divergence $d(\tau,\upsilon)$ between state visitation distributions of trajectories $\tau$ and $\upsilon$. Then, based on $d(\cdot,\cdot)$, we obtain the following trajectory-wise state marginal matching objective between $\pi$ and $\pi_{b}$:
@@
Compared with previous studies of state marginal matching [lee2019efficient, ghasemipour2020divergence], our approach emphasizes trajectory-level alignment. During training, many state-action pairs are concentrated near the starting point, and the agent only visits novel regions on rare occasions. Hence, due to the low data proportion, these novel state-action pairs have little impact on policy optimization when the SMM objective only considers the divergence between different state visitation distributions. Our approach addresses this problem by introducing a trajectory-wise SMM objective, which encourages the agent to engage in deep exploration.
Now the focus of the problem comes to the definition of $d(\cdot,\cdot)$. Inspired by previous research [masood2019diversity, Wang2023LearningDP], a trajectory can be considered a deterministic policy, and the distance between them can be defined as the MMD between their state visitation distributions. Therefore, $d(\cdot,\cdot)$ can be expressed as:
@@
$+\underset{s_{\upsilon},s_{\upsilon}^{\prime}\sim\rho_{\upsilon}}{\mathbb{E}}\left[k(s_{\upsilon},s_{\upsilon}^{\prime})\right]$
where $\rho_{\tau}$ and $\rho_{\upsilon}$ are the state visitation distributions of $\tau$ and $\upsilon$. To guarantee the stability of policy optimization, we further adopt the trust-region-based optimization method similar to TRPO [schulman2015trust]:
@@
$\text{s.t.}\ D_{\rm KL}(\pi_{\theta},\pi_{old})\leq\delta.$
Intuitively, this preference-guided objective function aids learning by orienting the agent’s policy toward human feedback. Specifically, our approach finds a new policy $\pi_{new}$ that minimizes the expected trajectory-wise MMD distance to $\pi_{b}$, while ensuring $\pi_{new}$ lies inside the trust region around $\pi_{old}$. This preference-guided policy optimization is the key distinguishing LOPE from other PbRL approaches. This method equips LOPE with two unique advantages over other state-of-the-art algorithms.
First, many previous PbRL studies explicitly learn separate reward functions and then employ off-the-shelf RL algorithms to train a control policy. In contrast, LOPE considers preference-labeled trajectories as guidance and reformulates a novel constrained optimization problem, as shown in Eq. (6). In this manner, LOPE effectively avoids the potential information bottleneck that occurs when conveying information from preferences to the policy via scalar rewards.
Second, LOPE learns a policy matching the trajectory-wise state visitation distribution of preference-labeled experiences. Consequently, LOPE enables the agent to quickly expand the scope of exploration along the preference-labeled trajectories in $\mathcal{P}$ and overcome the challenging exploration difficulty in online RL settings. Fundamentally, the proposed approach enables us to explore the available preference data more comprehensively, including position information, thereby providing more significant assistance to RL algorithms.